OpenStudy (anonymous):

One number is three times another number. Their sum is 52. Find the solution set. help on finding a way to solve this. Putting this into an equation would help. the fact that there are two variables is screwing with my head. so...help? =)

OpenStudy (binary3i):

let two numbers be x and y so x + y = 52 but x = 3y so 3y + y = 52 4y = 52 y = 13 x= 3*13= 39

OpenStudy (anonymous):

let's say first number is a and second number is b a = 3b a + b = 52 3b + b = 52 4b = 52 b = 13 a = 3*13 = 39

OpenStudy (anonymous):

thanks both of you ^_^ it really helped!

OpenStudy (anonymous):

No prob :D